You just purchased a house and have signed a 30-year mortgage for $500,000 with a 12% annual interest rate. What will be your monthly mortgage payment for this home?
Answer #1
Monthly payment = Loan amount / Present value of annuity of 1
= $       5,00,000 / 97.21833
= $       5,143.06
Working:
Present value of annuity of 1 = (1-(1+i)^-n)/i Where,
= (1-(1+0.01)^-360)/0.01 i = 12%/12 = 0.01
= 97.21833108 n = 30*12 = 360